\begin{thebibliography}{BMSS08}

\bibitem[BMSS08]{BMSS}
Alin Bostan, Fran{\c{c}}ois Morain, Bruno Salvy, and {\'E}ric Schost.
\newblock Fast algorithms for computing isogenies between elliptic curves.
\newblock {\em Mathematics of Computation}, 77(263):1755--1778, 2008.

\bibitem[DS05]{DiamondShurman}
Fred Diamond and Jerry~Michael Shurman.
\newblock {\em A First Course in Modular Forms}, volume 228.
\newblock Springer, 2005.

\bibitem[Elk98]{Elkies}
Noam~D. Elkies.
\newblock Elliptic and modular curves over finite fields and related
  computational issues.
\newblock In {\em Computational perspectives on number theory ({C}hicago, {IL},
  1995)}, volume~7 of {\em AMS/IP Stud. Adv. Math.}, pages 21--76. Amer. Math.
  Soc., Providence, RI, 1998.

\bibitem[Ful89]{Fulton}
William Fulton.
\newblock {\em Algebraic curves}.
\newblock Advanced Book Classics. Addison-Wesley Publishing Company, Advanced
  Book Program, Redwood City, CA, 1989.
\newblock An introduction to algebraic geometry, Notes written with the
  collaboration of Richard Weiss, Reprint of 1969 original.

\bibitem[Gal12]{Galbraith}
Steven~D. Galbraith.
\newblock {\em Mathematics of public key cryptography}.
\newblock Cambridge University Press, Cambridge, 2012.

\bibitem[Koh96]{Koh96}
David Kohel.
\newblock {\em Endomorphism rings of elliptic curves over finite fields}.
\newblock PhD thesis, University of California, Berkeley, 1996.

\bibitem[Sch95]{Schoof}
Ren{\'e} Schoof.
\newblock Counting points on elliptic curves over finite fields.
\newblock {\em Journal de th{\'e}orie des nombres de Bordeaux}, 7(1):219--254,
  1995.

\bibitem[Sil09]{Silverman}
Joseph~H Silverman.
\newblock {\em The Arithmetic of Elliptic Curves}, volume 106.
\newblock Springer, 2009.

\bibitem[Sta73]{Stark73}
H.~M. Stark.
\newblock Class-numbers of complex quadratic fields.
\newblock In {\em Modular functions of one variable, {I} ({P}roc. {I}nternat.
  {S}ummer {S}chool, {U}niv. {A}ntwerp, {A}ntwerp, 1972)}, volume Vol. 320 of
  {\em Lecture Notes in Math.}, pages 153--174. Springer, Berlin-New York,
  1973.

\bibitem[Sut13]{SutherlandEvaluation}
Andrew~V. Sutherland.
\newblock On the evaluation of modular polynomials.
\newblock In {\em A{NTS} {X}---{P}roceedings of the {T}enth {A}lgorithmic
  {N}umber {T}heory {S}ymposium}, volume~1 of {\em Open Book Ser.}, pages
  531--555. Math. Sci. Publ., Berkeley, CA, 2013.

\bibitem[{The}23]{sage}
{The Sage Developers}.
\newblock {\em {S}ageMath, the {S}age {M}athematics {S}oftware {S}ystem
  ({V}ersion 10.2)}, 2023.
\newblock {\tt https://www.sagemath.org}.

\bibitem[V\'71]{velu71}
Jacques V\'{e}lu.
\newblock Isog\'{e}nies entre courbes elliptiques.
\newblock {\em C. R. Acad. Sci. Paris S\'{e}r. A-B}, 273:A238--A241, 1971.

\end{thebibliography}

\documentclass{amsart}
\usepackage[utf8]{inputenc}

\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage{mathtools}

\usepackage{commath, graphicx}

\usepackage{amsmath, amsfonts, amsthm, amssymb, amscd, multirow, tikz}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{example}[theorem]{Example}

\theoremstyle{remark}

\newcommand{\cc}{\bb{C}}

\renewcommand{\S}{\sum}

\newcommand{\e}{\varepsilon}
\newcommand{\bb}{\mathbb}
\newcommand{\mc}{\mathcal}

\newcommand{\f}{\mathfrak}

\newcommand{\llp}{\left(}
\newcommand{\rrp}{\right)}

\usepackage{bookmark}

\newcommand{\Deg}{\operatorname{Deg}}
\DeclareMathOperator{\Char}{\operatorname{char}}

\renewcommand{\del}{\partial}

\newcommand{\ol}{\overline}

\newcommand{\jt}{\widetilde{j}}
\DeclareMathOperator{\End}{End}
\DeclareMathOperator{\Gal}{Gal}

\usepackage{xcolor}

\title{Computing isogenies at singular points of the modular polynomial}
\author{William E. Mahaney and Travis Morrison}
\date{}

\begin{document}
\maketitle

\begin{abstract}
In this paper we present a method which, given a singular point $(j_1, j_2)$ on $Y_0(\ell)$ with $j_1, j_2 \neq 0, 1728$ and an elliptic curve $E$ with $j$-invariant ${j_1}$, returns an elliptic curve $\widetilde{E}$ with $j$-invariant ${j_2}$ that admits a normalized $\ell$-isogeny $\phi\colon E\to \widetilde{E}$. The isogeny $\phi$ can then be efficiently computed from $E$ and $\widetilde{E}$ using an algorithm of Bostan, Morain, Salvy, and Schost. 
\end{abstract}

\section{Introduction}
For an elliptic curve $E$ defined over a field $k$, a separable isogeny $\phi\colon E\to E'$ is determined (up to post-composition with an isomorphism) by its kernel. 
This result is effective: for example, V\'{e}lu~\cite{velu71} gives formulas for computing $\phi$ (as a rational map) given the points in $\ker\phi$, and Kohel~\cite[Section 2.4]{Koh96} gives formulas for computing these rational maps from the {\em kernel polynomial} of $\phi$, the monic polynomial which vanishes precisely at the $x$-coordinates of the nontrivial points in $\ker\phi$. The methods of V\'{e}lu and Kohel also give an equation for $E'$, the codomain of $\phi$. The resulting isogeny is always {\em normalized}, meaning the invariant differential $\omega'$ of $E'$ pulls back, under $\phi$, to the invariant differential $\omega$ of $E$. The isogeny $\phi$ will be defined over $k$ if $\ker\phi$ is $\Gal(\overline{k})$-stable, or equivalently, if the kernel polynomial of $\phi$ has coefficients in $k$. 

As mentioned above, a (description of) a cyclic subgroup $K$ of $E$ determines a unique normalized isogeny $E\to  E/K$. Alternatively, given $E'$ and $d$ such that there exists a cyclic, normalized $d$-isogeny $\phi \colon E \to E'$, this isogeny can be computed with Stark's algorithm~\cite{Stark73} and even more efficiently with an algorithm due to Bostan, Morain, Salvy, and Schost~\cite{BMSS}. Suppose now one only knows there exists a not necessarily normalized isogeny between $E$ and $E'$  of prime degree $\ell$, perhaps because one has found that $\Phi_{\ell}(j(E),j(E'))=0$ where $\Phi_{\ell}(X,Y)$ is the {\em modular equation}.
Provided that $\Char(k)=0$ or is sufficiently large 
relative to $\ell$, an algorithm of 
Elkies~\cite[Section 3]{Elkies} can compute first a 
curve $\widetilde{E}\cong E'$ such that a normalized 
$\ell$-isogeny $\phi\colon E\to \widetilde{E}$ exists, 
and then the kernel polynomial for $\phi$,  
provided that $(j(E),j(E'))$ is a non-singular point on 
$\Phi_{\ell}(X,Y)=0$ (see also~\cite[Section 7]{Schoof} for further details and~\cite[Section 25.2.1]
{Galbraith} for pseudocode). This algorithm is an 
important subroutine in the SEA algorithm for point-counting on elliptic curves over finite fields: for an ordinary curve $E$ defined over $\bb{F}_q$, one computes a kernel polynomial for primes $\ell$ which split in $\End(E)\otimes\bb{Q}$ in order to compute the trace of the Frobenius endomorphism of $E$ modulo $\ell$. 

Elkies' method for computing an equation for the isogenous curve as described in~\cite{Schoof} works as follows. Suppose $E:y^2=x^3+Ax+B$ is an elliptic curve over $\bb{C}$ which is $\ell$-isogenous to the curve $E':y^2=x^3+A'x+B'$. Then there exists $q_0=\exp(2\pi i \tau_0)\in \bb{C}$ such that $A=-G_4(q_0)/48$ and $B=G_6(q_0)/864$ and $j(E)=j(q_0)$ where $G_4$ and $G_6$ are the Eisenstein series of weights $4$ and $6$ respectively and $j(q)$ is the $j$-function (in their $q$-expansions). Moreover, if we set $\widetilde{j}(q)\coloneqq j(q^{\ell})$, then $\widetilde{j}(q_0)=j(E')$ and setting 
\[
\widetilde{A} = \frac{-\ell^4}{48} \cdot \frac{ \jt(q_0)'^2}{\jt(q_0) (\jt(q_0) - 1728) },\quad 
\widetilde{B} = \frac{-\ell^6}{864} \cdot \frac{\jt(q_0)'^3}{\jt(q_0)^2 (\jt(q_0) - 1728) }
\]
yields coefficients for a short Weierstrass equation for $E'$ such that the $\ell$-isogeny $\phi\colon E\to \widetilde{E}$ is normalized. Therefore one can compute $\widetilde{E}$ as soon as one obtains $\widetilde{j}'(q_0)$. Letting $\Phi(X,Y)\coloneqq\Phi_{\ell}(X,Y)$ denote the $\ell$-th modular polynomial parametrizing $\ell$-isogenous $j$-invariant pairs, we have that $\Phi_{\ell}(j(q),\widetilde{j}(q))$ is identically zero as a power series in $q$. Differentiating with respect to $q$ yields the identity 
\[
j'(q_0)\Phi_X(j(q_0),\widetilde{j}(q_0))+\ell \widetilde{j}'(q_0)\Phi_Y(j(q_0),\widetilde{j}(q_0))=0
\]
so, {\em assuming $(j(E),j(E'))$ is a nonsingular point on $\Phi(X,Y)=0$}, one can solve for $\widetilde{j}'$ and hence $\widetilde{E}$ in terms of the equation for $E$ and the $j$-invariant of $E'$. 

If $(j(E),j(E'))$ is a singular $k$-point on $Y_0(\ell): \Phi_{\ell}(X,Y)=0$, there are (at least) two $\ell$-isogenies $\phi,\psi\colon E\to E'$ with distinct kernels. 

Schoof writes that the approach to computing the isogenous curve $\widetilde{E}$ ``does not work if $\Phi_X(j,\widetilde{j})=\Phi_Y(j,\widetilde{j})=0$.'' However, the broader idea -- compute $\widetilde{j}'$ in order to find the model  $\widetilde{E}$ --  does. In this paper, we take up the problem of computing an $\ell$-isogeny $E\to E'$ when $(j(E),j(E'))$ is a singular point on $Y_0(\ell)$. One simply needs more derivatives of the identity $\Phi_{\ell}(j,\widetilde{j})=0$. 

In Section~\ref{sec:singular}, we find that if $(j_1,j_2)$ is a singular point of multiplicity $m$ on $\Phi_{\ell}(x,y)=0$, we can compute a degree $m$ polynomial whose roots yield the $m$ choices for $\widetilde{j}'$ and hence the $m$ models with $j$-invariant $j_2$ admitting normalized $\ell$-isogenies, under the assumption that $j(E),j(E')\not=0,1728$ (which is also assumed in~\cite{Schoof}) and the characteristic of the field is zero or is sufficiently large relative to $\ell$.  With the equation for $\widetilde{E}$, the algorithm of~\cite{BMSS} computes $\phi$ efficiently. This allows one to compute {\em all} $\ell$-isogenies of $E$ efficiently. 

In Section~\ref{sec:background}, we  recall relevant definitions and notation. In Section~\ref{sec:elkies}, we review Elkie's method for computing an equation for the isogenous curve. In Section~\ref{sec:singular}, we extend Elkies' method for computing the equation for the isogenous curve at a singular point of the modular equation. Finally, in Section~\ref{sec:algorithm}, we give pseudocode for our algorithm and an example. Our implementation in Sagemath~\cite{sage} is available at \url{https://github.com/wmahaney/ModularMultipoints}.

\section{Background}\label{sec:background}
In this section, we include some background elliptic curves, modular curves, and singularities. We refer the reader to ~\cite{Silverman,DiamondShurman,Fulton} for further background.
\subsection{Elliptic Curves}

Let $k$ be a field. An elliptic curve over $k$ is a smooth projective genus $1$ curve with a $k$-rational point. We will assume throughout that the characteristic of $k$ is not $2$ or $3$ and that our elliptic curves are given by a {\em short Weierstrass equation over $k$}, that is, an equation of the form $y^2 =  x^3 + Ax + B$ where $x,y$ are indeterminates and $A, B \in k$ satisfy $4A^3 + 27B^2  \neq 0$. 
Given an elliptic curve $E: y^2 = x^3 + Ax + B$ the {\em  invariant differential} of $E$ is defined as \cite[III.5]{Silverman}
\[
\omega_E := \frac{dx}{2y} \in \Omega_E
\]
where $\omega_E$ is regarded as a differential on $E$. The {\em  $j$-invariant} of $E$ is given by  \[
j(E) := \frac{-1728(4A)^3   }{ -16(4A^3 + 27B^2)  }   = \frac{4 \cdot 1728A^3}{4A^3 + 27B^2}
\]
Two elliptic curves $E, E'$ defined over $k$ are $\ol{k}$-isomorphic if and only if $j(E) = j(E')$.  Defining
\[
E_j: y^2=x^3+3j(1728-j)x+2j(1728-j)^2,\quad j\not=0,1728\]
\[
 E_0: y^2=x^3+1,\quad E_{1728}:y^2=x^3+x
\]
produces curves $E_j$ such that $j(E_j)=j$ for any $j\in k$.

Given a pair of elliptic curves $E,E'$ an {\em isogeny} \cite[III.4]{Silverman} from $E \to E'$ is a non-constant morphism $\phi\colon E \to E'$ of projective varieties such that $\phi(\infty_{E}) = \phi(\infty_{E'})$.
Any isogeny from $E \to E'$ induces a group homomorphism $\phi\colon E \llp \ol{k} \rrp \to E' \llp \ol{k} \rrp$ \cite[III.4.8]{Silverman}. The {\em degree} of an isogeny is its degree as a rational map. Every isogeny $\phi\colon E\to E'$ has a {\em dual isogeny} $\widehat{\phi}\colon E'\to E$ which satisfies $\widehat{\phi}\circ\phi = [\deg\phi]$, the multiplication-by-$\deg\phi$ map. 
Two elliptic curves $E, E'$ are {\em $n$-isogenous}, for $n \ge 1$, provided there is an isogeny $\phi\colon E \to E'$ of degree $n$. We say that two $j$-invariants $j_1, j_2 \in K$ are {\em $n$-isogenous} provided $E_{j_1}, E_{j_2}$ are $n$-isogenous. 

An isogeny $\phi\colon E \to E'$ between elliptic curves is {separable} 
if it is separable as a rational map. If $\phi$ is separable, then $\deg\phi =\#\ker\phi$. 
We say that a separable isogeny $\phi\colon E \to E'$  is {cyclic} provided its kernel is a cyclic subgroup of $E(\ol{k})$. Every isogeny $\phi$ factors as $\phi=\phi'\circ[n]\circ\pi^n$ where $\pi$ is the $p$-power Frobenius isogeny, $[n]$ is the multiplication-by-$n$ map, and $\phi'$ is separable and cyclic.
And every $n$-isogeny $\phi\colon E \to E'$ can be factored as a composition of prime degree isogenies, so we will focus primarily on separable isogenies of prime degree.  

\subsection{Modular polynomials, curves, and functions}

Let $\ell$ be a prime. The modular curve $X_0(\ell)$ is an algebraic curve defined over $\mathbb{Q}$ whose non-cuspidal points parametrize elliptic curves with choice of $\ell$-isogeny. 
The curve $X_0(\ell)$ is given by the {\em modular equation} $\Phi_{\ell}(X,Y)=0$ which defines a plane curve $Y_0(\ell)$; we have that $(j_1,j_2)$ is a point on $Y_0(\ell)$ if and only if there are elliptic curves $E_{j_1},E_{j_2}$ and a cyclic $\ell$-isogeny $\phi\colon E_{j_1}\to E_{j_2}$. The curve $X_0(\ell)$ is the nonsingular projective model of $Y_0(\ell)$. Thus if $E$ is an elliptic curve and $K$ is a cyclic subgroup of $E$ of order $\ell$, then the point $[E,K]$ on $X_0(\ell)$ maps to the point $(j(E),j(E/K))$ on $\Phi_{\ell}(x,y)=0$. 
The modular polynomial $\Phi_\ell(X,Y)$ is irreducible, symmetric, and of the form $X^\ell Y^\ell + X^{\ell +1}+ Y^{\ell + 1} + \ldots$ where $\ldots$ denotes terms of lower degree \cite[Section 6]{Schoof}. 

There are some particular modular forms we are interested in for this paper, which we will regard in their $q$-expansions. The first are the weights 4 and 6 normalized (constant coefficient one) Eisenstein series $ G_4(q)$ and $G_6(q)$ which are modular forms of weights $4$ and $6$ respectively. There is also the $j$-function given by \[
j(q) := 1728 \frac{G_4(q)^3}{G_4(q)^3 - G_6(q)^2}
\]
The $j$-function parametrizes $j$-invariants of elliptic curves over $\cc$; in particular for a given $z \in \cc$ there exists a $q_0$ with $j(q_0) = z$ and the elliptic curve over $\cc$ with short Weierstrass equation \[
E: y^2 = x^3 - \frac{G_4(q_0)}{48} + \frac{G_6(q_0)}{864}
\]
has $j$-invariant $j(q_0)$. 

\subsection{Singularities of Plane Curves}
Two $j$-invariants $(j_1, j_2) \in k^2$ are $\ell$-isogenous over $k$ if and only if $\Phi_\ell(j_1, j_2) = 0$, so the $k$-rational points of $Y_0(\ell)/k$ parameterize pairs of $k$-rational $\ell$-isogenous $j$-invariants. While $X_0(\ell)/ k$ is smooth, $Y_0(\ell)/k$ is not generally smooth. Before discussing what singularities $Y_0(\ell)/k$ has let us define singularities and their associated data:

\begin{definition} (Singularities of Plane Curves)
\cite[Section 3.1]{Fulton} \\
Let $F(x,y) \in k[x,y]$ be a non-constant polynomial. A point $(a,b) \in k^2$ is a {\em multiple point} or {\em singular point} of $F$ provided \[
F(a,b) = \frac{\del F}{\del x} (a,b) = \frac{\del F}{\del y}(a,b) = 0
\]
where the derivatives are formal. 
Suppose $(a,b)$ is a singular point of $F$. Via change of variables define the polynomial \[
H(x, y) = F(x+a, y+b).
\]
As $H(0,0) = F(a,b) = 0$ we can write \[
H(x,y) = H_m(x,y) + \ldots + H_{\Deg(H)}(x,y),
\]
where each $H_d$ is homogeneous of degree $d$ and $H_m \neq 0$. The {multiplicity} of the singularity $(a,b)$ of $F$ is $m$. 
For a curve $C = \mc{V}(F)$, a point $(a,b)$ on $C$ is a singular point provided $(a, b)$ is a singular point of $F$ and vice versa. 
\end{definition}

\section{Elkies' method to compute the isogenous curve}\label{sec:elkies}
In this section, we review ideas of Elkies~\cite[Section 3]{Elkies} and formulas of Schoof~\cite[Section 7]{Schoof} for computing, given an elliptic curve $E$ and prime $\ell$, an equation for an elliptic curve $\widetilde{E}$ such that there exists a normalized $\ell$-isogeny $\phi\colon E\to \widetilde{E}$. 

Per \cite[Proposition 7.1 (i)]{Schoof} we have an equality of power series \begin{align*}
\frac{j'(q)}{j(q)} = - \frac{G_6(q)}{G_4(q)}
\end{align*}
where the above functions are considered as modular forms in their $q$-expansions; this equality gives another \[
j'(q) = - \frac{G_6(q)}{G_4(q)} j(q)
\]
away from $q_0$ for which $j(q_0) \in \{0, \infty\}$.  
Per \cite[Proposition 7.2]{Schoof} there exist formal power series $X(\zeta; q), Y(\zeta; q) \in \bb{Z} \Big{[}\Big{[} q, \zeta, \frac{1}{\zeta - \zeta^2} \Big{]} \Big{ ] }$ for which \begin{align*}
Y(\zeta; q)^2 = X(\zeta; q)^3 - \frac{G_4(q)}{48} X(\zeta; q) + \frac{G_6(q)}{864} 
\end{align*}
If one took the substitution \begin{align*}
A(q) & := \frac{-G_4(q)}{48} \\
B(q) & := \frac{G_6(q)}{864}
\end{align*}
Then the prior equality of power series becomes \[
Y(\zeta; q)^2 = X(\zeta; q)^3 + A(q)X(\zeta; q) + B(q)
\] and each fixed $q_0$ we have a short Weierstrass equation for an elliptic curve over $\cc$ whose affine points are the pairs $\llp X(\zeta(z); q_0), \pm Y(\zeta(z); q_0) \rrp$ where $\zeta(z) = e^{2i \pi z}$ and $z$ ranges over $\cc$. 
Fix a particular $q_0$. Then the curve $E: Y^2 = X^3 + A(q_0)X + B(q_0)$ is $\ell$-isogenous to the curve $\widetilde{E}\colon Y^2 = X^3 + A(q_0^\ell) X + B(q_0^\ell)$, and if one replaces $A(q_0^\ell), B(q_0^\ell)$ with $\ell^4 A(q_0^\ell)$ and $\ell^6 B(q_0^\ell)$ then the isogeny connecting the two curves is normalized~\cite{Elkies}. Using the identities presented in \cite[Proposition 7.1]{Schoof} one can compute formulas for $G_4(q_0), G_6(q_0)$ in terms of $j(q_0)$ and $j'(q_0)$, similarly one can compute formulas for $G_4(q_0^\ell)$ and $G_6(q_0^\ell)$ in terms of $\jt(q_0)$ and $\jt'(q_0)$. Therefore if one can compute $\jt'(q_0)$ one can obtain a model for $\widetilde{E}$ which admits an isogeny from $E$, and by scaling this model appropriately compute a normalized model for this isogeny. 

Fix a prime $\ell$ and the function $\jt(q) = j(q^\ell)$. For every $q_0$ we have that $j(q_0), \jt(q_0)$ are $\ell$-isogenous $j$-invariants so as a function of $q$ we have $\Phi_\ell(j, \jt) = 0$. Moreover if we take the derivative of the function $0 = \Phi_{\ell}(j, \jt)$ by applying the multivariate chain rule we get 
\[
0 = \llp \Phi_{\ell}(j, \jt) \rrp' = j' \Phi_X(j, \jt) + \ell \jt' \Phi_Y(j, \jt).
\]
Solving for $\widetilde{j}'$, assuming $\Phi_y(j, \jt) \neq 0$, and then setting 
\begin{align}
\widetilde{A} = \frac{-\ell^4}{48} \cdot \frac{ \jt'^2}{\jt (\jt - 1728) } \\
\widetilde{B} = \frac{-\ell^6}{864} \cdot \frac{\jt'^3}{\jt^2 (\jt - 1728) }
\end{align}
yields an equation $Y^2=X^3+\widetilde{A}X+\widetilde{B}$ for the isogenous curve admitting a normalized isogeny from $E$ whose points are the pairs $\llp X(\zeta(z); q_0^\ell), \pm Y(\zeta(z); q_0^\ell) \rrp$ as $z$ ranges over $\cc$. 

Even though all the above is derived from properties of modular functions over $\cc$, the Deuring lifting theorem justifies that these computations can be performed over a finite field as well, provided the characteristic is not $2,3$, or $\ell$.

\section{Singular points on \texorpdfstring{$Y_0(\ell)$}{Y0(l)}}\label{sec:singular}
In this section, we observe that the multiplicity of a point $(j_1,j_2)$ on $Y_0(\ell)$ is equal to the number of cyclic $\ell$-isogenies from an elliptic curve with $j$-invariant $j_1$ to an elliptic curve with $j$-invariant $j_2$. We use this in the following section in order to compute normalized models for $\ell$-isogenies at singular points of $Y_0(\ell)$.

\begin{proposition}
Suppose $j_1,j_2$ are two elements of $k$, neither of which are $0$ or $1728$. Let $(j_1,j_2)$ be a point on $Y_0(\ell)$ and let $m$ be the number of  $\ell$-isogenies with distinct kernels between two fixed elliptic curves $E_1$ and $E_2$ with $j$-invariants $j_1$ and $j_2$. If $\Char(k)=0$ or $\Char(k)>m$, then $(j_1,j_2)$ is a point of multiplicity $m$ on $Y_0(\ell)$. In particular,  all partial derivatives of $\Phi_\ell(X,Y)$ of weight less than $m$ vanish at $(j_1,j_2)$. Moreover,  both $\Phi_{X^m}(j_1,j_2))$ and $\Phi_{Y^m}(j_1,j_2)$ are nonzero.    
\end{proposition} \label{prop:partialsvanish}

\begin{proof}
We can identify $X_0(\ell)$ as the blowup of $Y_0(\ell)$ at its singular points. The non-cuspidal points on $X_0(\ell)$ then correspond to tangent directions at points $(j_1,j_2)$ of $Y_0(\ell)$. Suppose there are $m$ distinct points on $X_0(\ell)$ above $(j_1,j_2)$. Then the multiplicity of $(j_1,j_2)$ on $Y_0(\ell)$ is at least the number of tangent directions at $P$ and therefore at least $m$. On the other hand, the number of $\ell$-isogenies between $j_1$ and $j_2$ is equal to the multiplicity of $j_2$ as a root of the degree $\ell+1$ polynomial $\phi(Y)\coloneqq \Phi(j_1,Y)$. If $m<\ell+1$, the assumption on the characteristic of $\Char(k)$ implies $\phi^{(m)}(j_2)\not=0$ and therefore the multiplicity is at most $m$. If $m=\ell+1$ then $(j_1,j_2)$ is a point of multiplicity $m$ since its multiplicity is at most $\ell+1$. 
\end{proof}

In the following section, we will need to compute $\widetilde{j}'$ at singular points by finding it as a root of a polynomial obtained by differentiating the modular equation $m+1$ times, where $m$ is the multiplicity of the singularity. A straightforward argument via induction gives the following expression for the $n$th derivative of the modular equation:
\begin{proposition} \label{prop:diffeq}
Fix the functions $j(q)$ and $\widetilde{j}(q) \coloneqq j(q^\ell)$ in their $q$-expansions as well as the $\ell$th modular polynomial $\Phi = \Phi_\ell(X,Y)$. 
Then for all integers $n \ge 2$ the $n$-fold derivative of $\Phi(j, \jt)$ is given by \[
\llp \Phi(j, \jt) \rrp^{(n)} = \e_n(q) + F_n(q)
\]
where $\e_n(q)$ is of the form
\[
\e_n(q) = \S_{0 \le u+v < n} \e_{u,v}(q) \Phi_{X^u Y^v}(j, \jt)
\]
and $F_n(q)$ is of the form \[
F_n(q) = \S_{u=0}^n \binom{n}{u} \ell^{n-u} (j')^u (\jt')^{n-u} \Phi_{X^u Y^{n-u}}(j, \jt).
\]
\end{proposition}

\section{Computing Reductions of \texorpdfstring{$\widetilde{j}'$}{j~'}}
Throughout this section, fix a field $k$, fix $j_1,j_2\in k$ with $j_1,j_2\not=0,1728$, let $E_{A,B}$ be a curve of $j$-invariant $j_1$, and define $j_1' = \frac{18B}{A}j_1$. 


\begin{definition} \label{def:FiberPolynomial}
    Suppose $P=(j_1,j_2)$ is a singular point of multiplicity $m$ on $Y_0(\ell)$ with $j_1,j_2\not=0,1728$. Define 
    \[
F_P(t)= \S_{u=0}^m \binom{m}{u} \ell^{m-u} (j_1')^u\Phi_{X^u Y^{m-u}}(P) t^{m-u}.
\]
\end{definition}
\begin{remark}
By Proposition~\ref{prop:partialsvanish} $F_P(t)$ is a degree $m$ polynomial whenever $m < \Char(k)$. 
\end{remark}

\begin{proposition} \label{prop:models_are_roots}
Fix a singular point $P = (j_1, j_2)$ of multiplicity $m$ on $Y_0(\ell)/k$ and the associated polynomial $F_P(t)$. Then for every $k$-model $(\widetilde{A}, \widetilde{B})$ for $E_{j_2}$ admitting a normalized cyclic $\ell$-isogeny from $E_{A, B}$ there exists a root $\widetilde{r}$ of $F_P(t)$ given by $\widetilde{r} = \frac{18 \widetilde{B}}{\ell^2 \widetilde{A} }j_2$.
\end{proposition}

\begin{proof}
    Let $\phi\colon E_{A,B}\to E_{ \widetilde{A}, \widetilde{B}}$ be a normalized cyclic $\ell$-isogeny defined over $k$ and define 
    \[
    \widetilde{r} = \frac{18\widetilde{B}}{\ell^2 \widetilde{A}}j_2.
    \]
Let $\varphi\colon \widehat{E_{A,B}}\to \widehat{E_{\widetilde{A},\widetilde{B}}}$ be a lift of $\phi$ to an $\ell$-isogeny of elliptic curves defined over a number field $L$ such that there exists a prime $\f{P}$ of $L$ where $\widehat{E_{A,B}}$ and $\widehat{E_{\widetilde{A},\widetilde{B}}}$ both have good reduction at $\f{P}$ and $\varphi$ reduces to $\phi$ modulo $\f{P}$; in the case that $k$ is a number field we take $L = k$ and $\f{P}$ is the zero ideal. There exists $q_0$ such that 
\begin{align*}
\frac{- G_4(q_0)}{48} & \equiv A\pmod{\f{P}} & \frac{G_6(q_0)}{864} &\equiv B \pmod{\f{P}} \\ 
\frac{-\ell^4 G_4(q_0^\ell)}{48} & \equiv \widetilde{A} \pmod{\f{P}} & \frac{\ell^6 G_6(q_0^\ell)}{864} & \equiv \widetilde{B} \pmod{\f{P}} 
\end{align*}
and we can realize $\varphi$ as the isogeny mapping \[
\widehat{E_{A, B}}: y^2 = x^3 - \frac{G_4(q_0)}{48} x+ \frac{G_6(q_0)}{864}
\]
to
\[
\widehat{E_{\widetilde{A}, \widetilde{B}}}: y^2 = x^3 - \frac{\ell^4 G_4(q_0)}{48}x + \frac{\ell^6 G_6(q_0)}{864}
\]
which is normalized over $L$~\cite[Section 3]{Elkies}. 
We then have an equality of power series giving an equivalence \[
\frac{\jt'(q_0)}{\jt(q_0)} = - \frac{ \ell^6 G_6(q_0^\ell)}{\ell^4 G_4(q_0^\ell)} \equiv \frac{-864 \ell^4 \widetilde{B}}{-48 \ell^6 \widetilde{A}} \equiv \frac{18 \widetilde{B}}{\ell^2 \widetilde{A}} \pmod{\f{P}}.
\]
Because $0 \neq j_2 \equiv \jt(q_0) \bmod{\f{P}}$ we have $\jt(q_0) \neq 0$ and so \[
\jt'(q_0) \equiv \frac{ 18 \widetilde{B} \jt(q_0) }{ \ell^2 \widetilde{A}} \equiv  \frac{18 \widetilde{B}}{\ell^2 \widetilde{A}} j_2 \equiv \widetilde{r} \pmod{\f{P}}
\]
So $\widetilde{r}$ is a reduction of $\jt'$ modulo $\f{P}$. As $j(q), \jt(q)$ are $j$-invariants for a pair of $\ell$-isogenous elliptic curves over $\cc$ for all $q$ we have that \[
 \Phi_{\ell}(j, \jt) = 0
\]
Hence the $m$th derivative $\llp \Phi_{\ell}(j, \jt) \rrp^{(m)} = 0$ as well. Using Proposition~\ref{prop:diffeq} and the prior observation we can write
\[
0 = \llp \Phi(j, \jt) \rrp^{(m)}  = \e_m(q) + \S_{u=0}^m \binom{m}{u} \ell^{m-u} (j')^u (\jt')^{n-u} \Phi_{X^u Y^{m-u}}(j, \jt)
\]
where
\[
\e_m(q) = \S_{0 \le u+v < m} \e_{u,v}(q) \Phi_{X^u Y^v}(j, \jt)
\]
As $(j_1, j_2)$ is a singular point of $Y_0(\ell)/k$ of multiplicity $m$ by ~\ref{prop:partialsvanish} we have $\Phi_{X^u Y^v}(j_1, j_2) = 0$ whenever $u+v<m$. As $j_1 \equiv j(q_0) \pmod{\f{P}}$ and $j_2 \equiv \jt(q_0) \pmod{\f{P}}$ we have \begin{align*}
\e_m(q_0) & = \S_{0 \le u+v < m} \e_{u,v}(q_0) \Phi_{x^u y^v}(j(q_0), \jt(q_0)) \\
& \equiv \S_{0 \le u+v < m} \e_{u,v}(q_0) \Phi_{x^u y^v}(j_1, j_2) \pmod{\f{P}} \\
& \equiv 0 \pmod{\f{P}}
\end{align*}
So both $\llp \Phi(j, \jt) \rrp^{(m)}(q_0)$ and $\e_m(q_0)$ are zero modulo $\f{P}$, hence their difference is also zero modulo $\f{P}$ hence \begin{align*}
0 & \equiv H^{(m)}(q_0) - \e_m(q_0) \pmod{\f{P}} \\
& \equiv \S_{u=0}^m \binom{m}{u} \ell^{m-u} (j'(q_0))^u (\jt'(q_0))^{n-u} \Phi_{X^u Y^{m-u}}(j(q), \jt(q)) \pmod{\f{P}}  \\
& \equiv \S_{u=0}^m \binom{m}{u} \ell^{m-u} (j_1')^u (\widetilde{r})^{n-u} \Phi_{X^u Y^{m-u}}(j_1,j_2)  \pmod{\f{P}}\\
& \equiv F_P(\widetilde{r}) \pmod{\f{P}}
\end{align*}
\end{proof}
Thus we have a map from the set of normalized isogenies $\phi\colon E_{A, B} \to E_{\widetilde{A}, \widetilde{B}}$ to the set of roots of $F_P(t)$. As $P$ is a singular point of multiplicity $m$ there exist $m$ distinct cyclic $\ell$-isogenies from $j_1$ to $j_2$. However it can happen that there are two (or more) normalized $\ell$-isogenies between two fixed elliptic curves. 
\begin{example}
     Consider the curves $E:y^2=x^3+x+4$ and $\widetilde{E}:y^2=x^3+12x+7$ over $\bb{F}_{169}$. These curves are isomorphic, supersingular, and are connected by two normalized $\bb{F}_{169}$-rational $5$-isogenies with distinct kernels. The kernel polynomials are the two degree $2$ factors of $x^4 + 10x^3 + 9x^2 + 12x + 6$ over $\bb{F}_{169}$. 
\end{example}

\begin{proposition}\label{prop:isogeny_FP_root_bijection}
    Assume now that either $\Char(k)=0$ or if $\Char(k)=p>0$ that $p>4\ell$. Then there is a bijection between the roots of $F_P(t)$ and the cyclic, normalized $\ell$-isogenies of $E_{A,B}$. 
    Provided the bijection exists it is given by the following: 
    \begin{enumerate}
    \item Given a model $(\widetilde{A},  \widetilde{B})$ for $E_{j_2}$ admitting a normalized cyclic separable $\ell$-isogeny from $E_{A, B}$ the quantity $r = \frac{18 \widetilde{B}}{\ell^2 \widetilde{A}} j_2$ is a root of $F_P(t)$.
\item Given a root $r$ of $F_P(t)$ if one sets \[
   \widetilde{A} = \frac{- \ell^4}{48} \frac{r^2}{j_2(j_2-1728)}, \quad \widetilde{B} = \frac{-\ell^6}{864}\frac{r^3}{j_2^2(j_2-1728)},
\]
then $E_{\widetilde{A}, \widetilde{B}}$ admits a normalized cyclic separable $\ell$-isogeny from $E_{A, B}$. 
\end{enumerate}
\end{proposition}
For this, we need the following lemma:
\begin{lemma}\label{lem:multi_normalized}
    Let  $E,E'$ be elliptic curves over a field $k$ and whose $j$-invariants are not $0$ or $1728$.  Suppose there exist two normalized, separable $\ell$-isogenies $\phi,\psi\colon E\to E'$ with distinct kernels. Then $0<\Char(k)<4\ell$. 
\end{lemma}
\begin{proof}
The existence of two distinct subgroups of order $\ell$ in $E$ implies that $E[\ell]\cong (\bb{Z}/\ell\bb{Z})^2$ and in particular $\Char(k)\not=\ell$. Consider the (necessarily separable) endomorphism $\alpha=\widehat{\psi}\circ\phi\in\End(E)$. Since $\ker\phi\not=\ker\psi$, the endomorphism $\alpha$ is cyclic of degree $\ell^2$ and therefore $\alpha\not=\pm \ell$. Since both $\phi$ and $\psi$ are normalized, we have that $\alpha^*\omega=\ell\omega$ for any invariant differential of $E$. Therefore the endomorphism $\alpha-\ell$ is inseparable, so $p=\Char(k)>0$ and the degree of $\alpha-\ell$ is divisible by $p$. Let $t(\alpha)=\alpha+\widehat{\alpha}\in \mathbb{Z}$ denote the trace of $\alpha$.  The degree of $\alpha-\ell$ is  
\[
\deg(\alpha-\ell) = 2\ell^2-\ell t(\alpha) \equiv 0\pmod{p},
\]
Since $\alpha$ is separable, the prime $p$ does not divide $\ell$, so $t(\alpha)-2\ell\equiv 0 \pmod{p}$.   On the other hand, we have that the discriminant $t(\alpha)^2-4\ell^2$ is negative, so $-2\ell<t(\alpha)<2\ell$. This implies 
\[
-4\ell<t(\alpha)-2\ell<0
\]
so $p<4\ell$, as claimed.
\end{proof}
We are now ready to prove Proposition~\ref{prop:isogeny_FP_root_bijection}.
\begin{proof}[Proof of Proposition~\ref{prop:isogeny_FP_root_bijection}]

    Since $P$ is a point of multiplicity $m$ on $Y_0(\ell)$, there exist $m$ different $\ell$-isogenies of $E=E_{A,B}$ with distinct kernels to curves with $j$-invariant $j_2$. Let $(\widetilde{A_1},\widetilde{B_1}),\ldots,(\widetilde{A_m},\widetilde{B_m})$ denote the coefficients of the short Weierstrass equations for these $m$ codomains. By Lemma~\ref{lem:multi_normalized}, these curves are all distinct. On the other hand, since they are all isomorphic, for each $i,j$ there exists $u_{ij}\in \overline{k}$ such that $\widetilde{A_i} = u_{ij}^4\widetilde{A_j}$ and $\widetilde{B_i}=u_{ij}^6\widetilde{B_j}$. Define $r_i = \frac{18}{\ell^2}{j_2}\frac{\widetilde{B_i}}{\widetilde{A_i}}$ for each $1 \le i \le m$. We claim that the $r_i$ are distinct. 
    
      Observe that if 
    \[
    r_i = {j_2}\frac{18\widetilde{B_i}}{\ell^2\widetilde{A_i}} = {j_2}\frac{18\widetilde{B_j}}{\ell^2\widetilde{A_j}} = r_j
    \]
    then 
    \[
     \frac{\widetilde{B_j}}{\widetilde{A_j}} = \frac{\widetilde{B_i}}{\widetilde{A_i}} =u_{ij}^2 \frac{\widetilde{B_j}}{\widetilde{A_j}}
    \]
    so $u_{ij}=\pm1$, which would imply $(\widetilde{A_i},\widetilde{B_i})=(\widetilde{A_j},\widetilde{B_j})$. Thus $r_i\not=r_j$ for $i\not=j$. By Proposition ~\ref{prop:models_are_roots} each of the $r_i$ is a root of $F_P$, thus we have constructed $m$ distinct roots of the degree $m$ polynomial $F_P$ hence as claimed each root of $F_P$ corresponds to a model $(\widetilde{A}_n, \widetilde{B}_n)$ and vice versa. 
    The first formula to obtain $r_i$ from $\widetilde{A}_i, \widetilde{B}_i$ is proven in ~\ref{prop:models_are_roots}. The formula to obtain $\widetilde{A}_i, \widetilde{B}_i$ from $r_i$ is proven by taking Deuring lifts of $\widetilde{A}_i, \widetilde{B}_i, j_2$ and $\jt'$ and expressing these lifts in terms of Eisenstein series, then one uses the power series identities in \cite[Proposition 7.1]{Schoof} applied to $j(q_0^\ell), \jt(q_0)^\ell, G_4(q_0^\ell), G_6(q_0^\ell)$ to manipulate the expressions for $r_i^2, r_i^3$ and obtain the claimed formulas.

\end{proof}

\section{Pseudocode and examples}\label{sec:algorithm}

~\ref{prop:isogeny_FP_root_bijection} gives a polynomial time algorithm which upon input $(j_1, j_2)$ a multiplicity $m$ point on $Y_0(\ell)$ returns every model for $E_{j_2}$ which permits a normalized isogeny from $j_1$. We simply compute the polynomial given in definition ~\ref{def:FiberPolynomial}, find its roots, and compute the $\widetilde{A}, \widetilde{B}$ corresponding to each root we computed. 
\begin{algorithm}[ht!]
\caption{Basic Algorithm To Compute Normalized Models for the Isogenous Curve over Finite Fields}\label{alg:cap}
\begin{algorithmic}
\Require A prime $\ell \neq \Char(k)$ such that $4 \ell \le \Char(k)$ or $\Char(k) = 0$, a singular point $(j_1, j_2) \in \mc{V}_k(\Phi_{\ell}(X, Y))$ of multiplicity $m$ with $j_1,j_2 \neq 0,1728$, and a $k$-model $(A, B)$ for $E_{j_1}$. 
\Ensure $\{ (A_1, B_1), \ldots, (A_m, B_m) \}$ such that for all $1 \le i \le m$ the cyclic $\ell$-isogeny from $E_{j_1}$ with model $(A,B)$ to $E_{j_2}$ with model $(A_i, B_i)$ is normalized.
\State $\Phi \gets \Phi_{\ell}(X,Y) \in k[X,Y]$ \Comment{Compute the $\ell$th modular polynomial.} 
\State $j_1' \gets \frac{18B}{A}j_1$  \Comment{Compute the reduction of $j_1':=j'(q_0)$ to $k$.}
\For{$0 \le u \le m$}
\State $c_u \gets \binom{m}{u} \ell^{m-u} (j_1')^u \Phi_{X^u Y^{m-u}}(j_1, j_2)$ \Comment{Compute Coefficients of $F_P$.}
\EndFor
\State $F(t) \gets \S_{u=0}^m c_u t^u$ \Comment{Instantiate $F_P$.}
\State $\{r_1, \ldots, r_m \} \gets \text{Roots{$(F)$}}$ \Comment{Compute the roots of $F_P$.}
\For{$1 \le i \le m$} \Comment{Each $r_i$ is a valid $j_2'$.} 
\State $A_i  = - \frac{\ell^4}{48} \frac{r_i^2}{j_2 (j_2 - 1728) } $
\State $B_i = - \frac{\ell^6}{864} \frac{r_i^3}{j_2^2(j_2 - 1728)}$
\EndFor
\State {Return} $\{ (A_1, B_1), \ldots, (A_m, B_m) \}$. 
\end{algorithmic}
\end{algorithm}

\begin{example}
Take $k = \mathbb{F}_{137^2} \cong \mathbb{F}_{137}(w)$ where $w$ has minimal polynomial $x^2 + 131x + 3$ over $\mathbb{F}_{137}$. Fix $j_1 = 136$ and $j_2 = 22$ regarded in $k$, as well as $P = (j_1, j_2) \in k^2$ and $\ell = 5$. Note that $1728 \equiv 84 \pmod{137}$ so neither $j_1$ nor $j_2$ are in $\{0, 1728\}$. 
The instantiated modular polynomial $\Phi_5(136,y)$ has a double factor of $y-22$ so there are two distinct 5-isogenies from $j_1 = 136$ to $j_2=22$. Fix a model $E_{A, B}: y^2 = x^3 + 19x + 65$ for $E_{136}$ and compute $j_1' = \frac{18 \cdot 65}{19} \cdot 136 \equiv 61$. We compute the partial derivative evaluations $\Phi_{XX}(P), \Phi_{XY}(P)$, and $\Phi_{YY}(P)$ as 
\begin{align*}
\Phi_{XX}(P)= 79, \quad \Phi_{XY}(P) = 6, \quad \Phi_{YY}(P)  = 5
\end{align*}
We can now compute $F_P(t)$ as \begin{align*}
F_P(t) & = \binom{2}{0} \ell^{2} (j_1')^0 \Phi_{YY}(P) t^2 + \binom{2}{1} \ell^{2-1} (j_1')^{1} \Phi_{XY}(P)t + \binom{2}{2} \ell^{2-2} (j_1')^2 \Phi_{XX}(P) t^0 \\ 
& = 5^2 \cdot 5 \cdot t^2 + 2 \cdot 5 \cdot 61 \cdot 6t + 61^2 \cdot 79 \\
& \equiv 125t^2 + 98t + 94 \pmod{137}
\end{align*}
The roots of $F_P(t)$ are \[
r_1 = 134w+93, \quad r_2 = 3w+75
\] 
and the associated models are given by \[
\widetilde{A}_1 = 32w+118, \quad \widetilde{B_1} = 15w+136 , \quad \widetilde{A}_2 = 105w+136, \quad \widetilde{B}_2 = 122w+89
\]
The 5-isogeny from $E_{A, B} \to E_{\widetilde{A}_1, \widetilde{B}_1}$ has kernel polynomial \[
f_1 = x^2 + (32w+84)x+(107w+66)
\]
and the 5-isogeny from $E_{A, B} \to E_{\widetilde{A}_2, \widetilde{B}_2}$ has kernel polynomial \[
f_2 = x^2 + (105w+2)x+(30w+23)
\]
\end{example}

\begin{remark}
    In Algorithm~\ref{alg:cap}, we compute the $\ell$th modular polynomial in order to compute the polynomial $F_{(j_1,j_2)}(t)$ to compute $\widetilde{j}'$. Alternatively, one could modify Sutherland's algorithm~\cite{SutherlandEvaluation} for computing the instantiated modular polynomial $\phi(Y)\coloneqq \Phi(j_1,Y)$ and derivative $\phi_X(Y)\coloneqq\Phi_X(j_1,Y)$ to  
    additionally compute $\phi_{X^{i}}(Y)=\Phi_{X^{i}}(j_1,Y)$ for $i=1,2,\ldots,m$ at a multiplicity $m$ singular point $(j_1,j_2)$ of $Y_0(\ell)$ and then compute the polynomial $F_{(j_1,j_2)}$ using these instantiated polynomials. 
\end{remark}
\begin{remark}
    Elkies' method computes not only the model $\widetilde{E}$ such that a normalized $\ell$-isogeny $E\to \widetilde{E}$ exists; by computing the second derivative $\widetilde{j}''$, one can compute $\sigma$, the $(\ell-3)/2$ coefficient of the kernel polynomial, i.e. the sum of the $x$-coordinates of the affine points in the kernel. From this coefficient, the rest can be determined (see for example~\cite[Algorithm 28]{Galbraith} and~\cite[Propositions 7.1, 7.2, 7.3]{Schoof}). Schoof computes $\widetilde{j}''$ by differentiating the modular equation twice; one should be able to solve for $\widetilde{j}''$ at a singular point of multiplicity $m$ by differentiating the modular equation $m+1$ times. Now one would have to match possible values for $\widetilde{j}''$ with the possible values for $\widetilde{j}'$. This could be advantageous, since the algorithm of~\cite{BMSS} is faster (in terms of field operations) by a factor of $\log \ell$ when $\sigma$ is known. 
\end{remark}

\section{Acknowledgements}
We would like to thank Leonardo Mihalcea for helpful discussions. The first author was supported by the Julian Chin Ph.D. Fellowship in Cybersecurity.  This research was funded in part by the Commonwealth of Virginia’s Commonwealth Cyber Initiative (CCI). CCI is an investment in the advancement of cyber R\&D, innovation, and workforce development. For more information about CCI, visit \url{www.cyberinitiative.org.}

\bibliographystyle{alpha}
\bibliography{bib.bib}

\end{document}